40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Fri Sep 12 00:40:23 2014, Jenne, Update, LSC, DRMI locking 
    Reply  Fri Sep 12 11:25:42 2014, Koji, Update, LSC, DRMI locking 
       Reply  Fri Sep 12 12:00:59 2014, ericq, Update, LSC, DRMI locking 
    Reply  Fri Sep 12 14:11:17 2014, ericq, Update, LSC, DRMI locking 7x
       Reply  Wed Sep 17 15:40:00 2014, ericq, Update, LSC, DRMI locking dcAngleDiff_srcL.pdfdrfpmiVertexSensing.zip
          Reply  Tue Sep 23 17:37:10 2014, ericq, Update, LSC, DRMI locking SRCcorrection.pdf
             Reply  Wed Sep 24 11:02:38 2014, manasa, Update, LSC, Moving SRM 
                Reply  Thu Sep 25 11:33:41 2014, Jenne, Update, LSC, POY alignment laser 
Message ID: 10498     Entry time: Fri Sep 12 00:40:23 2014     Reply to this: 10500   10502
Author: Jenne 
Type: Update 
Category: LSC 
Subject: DRMI locking 

 Tonight I worked on DRMI locking.  

I think the reason the May2014 DRMI recipe wasn't working for me is because I wasn't including the REFL11 -> SRCL element.  I had left it out because (a) I didn't think we should need it and (b) REFL11 is going through the CM board.  

Tonight, I flipped the switch on the CM screen so that OUT2 was seeing REFL11I, not REFLDC, so I had REFL11 in the usual place.  I reset the demod phase, since we had left it at zero for CM stuff. 

Setting demod phases for PRMI:

I locked PRMI on sideband, REFL 33 I&Q and drove PRM.  REFL55 was at 55deg, and I changed it to 33deg to minimize the peak in the Q-phase.  REFL11 was a 0deg, and I set it to 17deg.  I also checked the AS55 phase in the MICH-only case, and changed it from 14.75deg to 24.75 deg.

The May 2014 recipe (elog 9968) calls for adding 25 degrees to the REFL55 phase, so I put REFL55 at 58deg for DRMI locking.

After that, using the parameters in the May2014 recipe, the DRMI just locked.  Awesome!

I checked the demod phases with DRMI lock.  REFL11 stays at 17 degrees.  If I actuate the SRM, I get the largest peak in the I-phase of REFL55 with a phase of -143deg, but the acquisition is best with phase around 55deg.  [Note, as Q points out, I wonder if SRCL is mostly locked with REFL11I for some magical reason, which is why it didn't matter so much that I put a sign flip into REFL55...I wonder if fixing our macroscopic length offset in SRCL will fix this].  I also changed the REFL165 phase from -155.5deg to +145deg.

By looking at transfer functions at an excitation frequency, I expected that I should be able to hold SRCL and MICH on REFL165, with matrix elements -0.085 for REFL165I->SRCL and -0.23 for REFL165Q->MICH.  I was not able to acquire with these values, nor was I able to ramp the matrix elements while keeping lock.

So, I tried moving PRCL to REFL33I, which did work.  I used 1.245*REFL33I->PRCL, but left SRCL and MICH on REFL55 I&Q, with the REFL11I->SRCL element also there. This is where I started trying to get rid of the REFL11I element, but couldn't maintain lock most times, and could never acquire lock without it.

Next up, checking the MICH->SRCL coupling due to the output matrix.  I did as Koji did in elog 8816 , but first I copied the notches in FM10 of MICH over to PRCL and SRCL (old notch freqs were SRCL=566.1Hz, PRCL=675.1Hz, now they're all 475.1Hz).  I drove BS, and checked that the PRM element minimized the peak in REFL33I, the PRCL error signal.  I also added an SRM element to reduce the peak in REFL55I, the SRCL error signal.  I ended up with 0.5*BS, -0.284*PRM, -1.5*SRM for MICH drive, and unity in the PRM and SRM elements for PRCL and SRCL, respectively.

I measured the SRCL open loop gain, and the UGF was pretty low, so I increased the SRCL gain from 0.2 to 0.5 to make the UGF be around 70Hz.  I measured PRCL and MICH also, and they matched their references.

I worked a little bit on trying to remove REFL11 from the SRCL error signal, but didn't get anywhere.  I'm leaving the IFO to Q for the rest of the night.

To sum up, here is the set of parameters that worked for DRMI locking.  (These are saved as the template on the IFO Config screen.):

DEMOD PHASES:

REFL11:  17 deg

REFL33: 140.5 deg (not changed tonight)

REFL55: 58 deg  (58deg for DRMI, 33deg for PRMI)

REFL165: 145 deg

AS55: 24.75 deg

INPUT MATRIX

MICH = 0.15 * REFL55Q

PRCL = 1.245 * REFL33I

SRCL = -0.09 * REFL11I + 1.0 * REFL55I

DOF Triggers

MICH, PRCL, SRCL all on POP22I, 50:10

GAINS

MICH = 1.0

PRCL = -0.02

SRCL = 0.5

FM triggers 

MICH:  35:2, 2 sec delay, FM 2, 3, 6, 9

PRCL:  35:2, 0.5 sec delay, FM 2, 3, 6, 9

SRCL:  35:2, 5 sec delay, FM 3, 6, 9  (always lose lock trying to engage FM2).

OUTPUT MATRIX

MICH = 0.5 * BS +  (-0.284)*PRM + (-1.5)*SRM

PRCL = 1*PRM

SRCL = 1*SRM

ELOG V3.1.3-